OpenStudy (anonymous):

Find x if: a. 3x+7=20 b. -2x+2=20

OpenStudy (diamondboy):

try for make x feel lonely

OpenStudy (diamondboy):

for a) subtract 7 from both sides and for b) subtract 2 from both sides

OpenStudy (anonymous):

what do mean when you say to subtract from both sides?

OpenStudy (anonymous):

For A- 3x+7=20 (Subtract 7 From a side to get x by itself) -7 -7 3x=13 (Divide by 3 to get x by itself) 3 3 x= 4/1 x=4 So the answer for A Is \[x =4\]

OpenStudy (anonymous):

And for B. -2x+2=20 (Isolate x) -2 -2 -2x = 18 (Divide by -2 to get x by itself) x = -9 So the answer for B is \[x =-9\]

OpenStudy (anonymous):

\[Metal \And Fan Please =D\]

OpenStudy (anonymous):

<3 Tag me if you need anything else =D